### v1

# A Perspective on Foundation Models for the Electric Power Grid

Abstract
Foundation models (FMs) currently dominate news headlines. They employ advanced deep learning architectures to extract structural information autonomously from vast datasets through self-supervision. The resulting rich representations of complex systems and dynamics can be applied to many downstream applications. Therefore, FMs can find uses in electric power grids, challenged by the energy transition and climate change. In this paper, we call for the development of, and state why we believe in, the potential of FMs for electric grids. We highlight their strengths and weaknesses amidst the challenges of a changing grid. We argue that an FM learning from diverse grid data and topologies could unlock transformative capabilities, pioneering a new approach in leveraging AI to redefine how we manage complexity and uncertainty in the electric grid. Finally, we discuss a power grid FM concept, namely GridFM, based on graph neural networks and show how different downstream tasks benefit.
keywords:
Foundation Models , Data-Driven Power Grid Modeling , Energy Transition , AI-based Power Flow Simulation
[…]
[ibmusa]organization=IBM Research,addressline=T.J. Watson Research Center, city=Yorktown Heights, postcode=10598, state=NY, country=USA
[…]
[ibmzrl]organization=IBM Research - Europe,addressline=Säumerstrasse 4, city=Rüschlikon, postcode=8803, country=Switzerland
[…]
[eq]organization=Equilibrium Energy, Inc.,addressline=333 Kearny Street, city=San Francisco, postcode=94108, state=CA, country=USA
[…]
[nrel]organization=National Renewable Energy Laboratory, addressline=15013 Denver W. Parkway, city=Golden, postcode=80401, state=CO, country=USA
[…]
[eth]organization=Reliability and Risk Engineering Lab, Institute of Energy and Process Engineering, Department of Mechanical and Process Engineering,addressline=ETH Zürich, city=Zürich, postcode=8092, country=Switzerland
[…]
[epfl]organization=École polytechnique fédérale de Lausanne (EPFL),addressline=Rte Cantonale, city=Lausanne, postcode=1015, country=Switzerland
[…]
[anl]organization=Argonne National Laboratory,addressline=9700 S Cass Av, city=Lemont, postcode=60439, state=IL, country=USA
[…]
[doe]organization=Department of Energy,addressline=1000 Independence Ave., S.W., city=Washington, DC, postcode=20585, state=DC, country=USA
[…]
[inesc]organization=INESC TEC,addressline=Campus da FEUP, Rua Dr Roberto Frias, city=Porto, postcode=4200-465, country=Portugal
[…]
[harvard]organization=Harvard University,addressline=150 Western Ave, city=Allston, postcode=02134, state=MA, country=USA
[…]
[ibmdub]organization=IBM Research - Europe,addressline=IBM Technology Campus Damastown Industrial Park Mulhuddart Co, city=Dublin, postcode=D15HN66, country=Ireland
[…]
[hq]organization=Hydro-Québec,addressline=1800 Bd Lionel-Boulet, city=Varennes, postcode=J3X1S1, state=QC, country=Canada
[…]
[uchicago]organization=University of Chicago,addressline=5730 S Ellis Ave, city=Chicago, postcode=60637, state=IL, country=USA
[…]
[cu]organization=University of Colorado,addressline=$1$ Engineering Dr, city=Boulder, postcode=80309, state=CO, country=USA
[…]
[nyu]organization=New York University,addressline=370 Jay Str, city=Brooklyn, postcode=11201, state=NY, country=USA
[…]
[bra]organization=Masaryk University,addressline=Kotlarska, 2, city=Brno, postcode=62000, country=Czech Republic
footnotetext: § Lead contact: hehamann.hh@gmail.com or hendrikh@us.ibm.comfootnotetext: * These authors contributed equally to this work as major contributors.footnotetext: † These authors contributed equally to this work as significant contributors.footnotetext: ‡ These authors contributed equally to this work.
1 Introduction
Power grids are arguably the largest machines ever built by humanity, as recognized by the National Academy of Engineering [1]. Having survived the deregulation revolution at the dawn of the new century, this aging and ever-expanding system is about to face its biggest challenge. The energy transition, with its widespread electrification and massive deployment of renewable and distributed generation, is the main driver for this change, affecting both the demand and supply sides of the market [2, 3, 4]. And it is happening fast. Climate change, manifesting with more frequent high-intensity weather events, poses additional challenges to system planners and operators [5, 6, 7].
Figure 1: The energy transition, aging infrastructure, cybersecurity challenges, and climate change greatly increase complexity and uncertainty in operating, controlling, and planning the power grid. This is creating a widening gap between existing computational capabilities and the evolving needs of the electric power industry. AC PF denotes alternating current power flow; DC PF denotes direct current power flow.
In this context, it is clear that the operation, control, and planning of power systems will soon be pushed to their limits. As depicted in Figure 1, new computational methods and approaches are needed, capable of better tackling the challenges presented by increased uncertainty and complexity. Machine learning (ML) and artificial intelligence (AI) methods have shown promise for such purposes in a wide spectrum of industries, with significant breakthroughs in computer vision [8], natural language processing [9], and intelligent control [10]. Such approaches rely on statistical algorithms to learn from massive data, often enabling them to execute tasks (e.g., regression, classification, detection, segmentation, forecasting, generation) without human supervision.
Yet, despite significant progress in the application of AI/ML methods to power systems [11, 12, 13], wider adoption has been impeded, in particular, by a lack of readily available training data and the limited transferability of developed models to adjacent applications [14]. This is where, we argue, Foundation models (FM) have an important role to play. FMs are advanced AI models developed through self-supervised learning, most often based on transformer architectures, that generalize across various tasks after initial training on large datasets, enabling efficient adaptation to specific applications with minimal labeled data. Text-based FMs like GPT-4 [15], encoding deep relationships in text, have enabled a wide range of new applications. Similarly, we anticipate that a multi-modal FM for power grids, trained on the diverse data that characterizes the power grid, will enable a broad set of new use cases.
To enable this, we advocate in this article for developing multi-modal FMs for power grids: what we refer to as GridFMs. We explain why we believe in their potential, highlighting their anticipated strengths and weaknesses in light of the challenges posed by the evolving power grid. We develop a concrete picture of what we envision a GridFM to be and how we expect it to pioneer a new approach in leveraging AI to manage complexity and uncertainty in the power grid.
In the following, we first highlight, in Section 2, the looming power grid challenges associated with the energy transition, which are further aggravated by aging infrastructure, swelling cybersecurity, and climate change concerns. We show how, by increasing the complexity and uncertainty of the power grid, these drivers are creating a significant gap in computational capabilities, which we propose to overcome with FMs. We then review in Section 3 the strengths and weaknesses of FMs and propose in Section 4 an implementation of a GridFM that we argue will enable exciting, currently unattainable use cases as downstream tasks.
2 Looming challenges
As the energy transition accelerates, power grids face unprecedented challenges threatening their reliability and efficiency.
2.1 Distributed and renewable energy sources
While a single coal plant can generate hundreds of megawatts predictably and reliably, large numbers of weather-dependent renewable energy resources like wind and solar are required to produce the same power output. Current hourly and daily forecast methods still show up to 3% and up to 10% forecast error, respectively [16]. Additionally, distributed energy resources (DERs) like battery storage and electric vehicles are operated by independent power management systems and are driven by user patterns with intrinsic uncertainties. Dealing with the added uncertainty from reverse DER power flows, multipoint injections, and weather-induced fluctuations requires new voltage regulation methods, protection schemes, and advanced (optimal) power flow analysis. The numbers of controllable and dispatchable devices, as well as balancing nodes are expected to grow exponentially [17] with additional thousands to millions to be added in the near future. Managing the associated variability and uncertainty will require advanced stochastic optimization, scenario reduction, and sampling-based approaches, which are computationally more intensive than currently used deterministic methods, especially for real-time demand-supply balancing [18]. Furthermore, as the net load of DER-rich distribution branches seen by transmission operators becomes more variable and less predictable, effective grid management and planning will become impossible without effective co-simulation frameworks that integrate transmission and distribution analysis. Yet the computational cost of such simulations, at least with existing technologies, is extremely high.
2.2 Everything inverter
For the future power grid, inverters will play a crucial role in maintaining low cost at the grid’s edge. However, with increased inverter proliferation, established control schemes and grid stability are challenged, while a more deregulated environment is created. With more inverters connected, grid inertia – which stabilizes phase, voltage, and frequency during transient events and which is mostly tied to large synchronous generators – decreases [19], unless more grid-forming converters are introduced. As inverters are driven by software, they can quickly adjust power or disconnect, causing additional issues like sub-synchronous resonances and even trip large generators [20].
Despite the increasing adoption of standards [21], inverters remain hard to predict and model, unlike physics-driven synchronous generators. Quasi-static phasor simulations are used to assess optimal power flow. For grid stability assessment, however, computationally expensive electro-magnetic transient (EMT) simulations [22] are required. They run mostly on traditional central processing units (CPUs), which are flexible for many algorithms but inferior in compute performance to recent graphics processing units (GPUs), which are just starting to get adopted in the industry, as, e.g., in [23]. Also, CPU systems do not always scale well for large EMT simulations [24]. Grid operations, however, would require EMT simulations to run in real-time and for large sections of the grid, including distribution grid sections. To further account for inherent network uncertainties and those introduced by inverter-based resources (IBRs), stochastic ensembles of the problems would need to be simulated. Altogether, such an approach is computationally expensive and, with current approaches, impractical [25]. Thus, the community mostly relies on offline simulators like PS CAD [26] or ParaEMT [27] while also exploring physics-informed neural networks to reduce EMT simulation time [28].
2.3 Changes to demand and weather patterns
Mass electrification of heating, transportation, and industry – in addition to climate change and shifting demographics – result in changing load profiles. Consequently, the accuracy of classical short-term load forecasting has declined over the past decade due, among other factors, to the changes in customer demand and weather patterns, leading to out-of-distribution scenarios [29]. To model demand uncertainty, utilities use multiple scenarios for structural and random drivers [30]. Increasingly, they also forecast at the power transformer level to manage local dynamics [31], but this approach is costly and does not scale due to the need for multiple models.
Short-term models
Short-term models for operations, running hourly to weekly, are used for market bidding, unit commitment/economic dispatch, operational planning, and security analysis, among others [32]. As stated before, uncertainty and variability are driven in part by weather and human behavior, from societal aggregations to individual households [33], as well as by operator-induced industrial activities. Economic factors, the increasing frequency and intensity of extreme weather events with high prediction uncertainty [34], and growing system complexity (e.g., self-consumption, electric vehicles) also contribute to variability. The resulting complexity necessitates the selection and integration of new and heterogeneous data sources (e.g., social networks, online news) along with the frequent adjustment, rebuilding, and maintenance of models.
Long-term models
Long-term models like the Regional Energy Deployment System [35] are used for resource, infrastructure, and strategic planning. Uncertainty and variability at longer time scales are significantly driven by climate policy, technology adoption, and regulatory changes. Instances of this include the combustion engine phase-out by 2035 in Europe – which resulted in a massive increase in the long-term electricity demand forecast – or the US Federal Energy Regulatory Commission’s rules on planning and covering the cost of the power grid [36]. To improve model accuracy, additional information that captures demographic changes and changes in the distribution of wealth, which lead to different societal and individual behaviors (e.g., heating, cooling, electromobility), as well as climate change-induced environmental shifts must be integrated. Therefore, the capacity to combine heterogeneous data sources (including information in textual format) is a fundamental requirement to improve predictability.
2.4 Aging infrastructure
Managing the vast amount of diverse equipment and aging infrastructure – the US power grid has hundreds of thousands of miles of transmission and millions of miles of distribution lines with over 55,000 substations – constitutes a major challenge on its own. While some lines last up to 80–100 years [37], insulators, transformers, and generators require frequent inspections, ideally through permanent sensing, which, given their age and approaching end-of-life [38], may not always be installed. Modeling these components in mostly unknown conditions is difficult and may rely on computationally expensive stochastic and ensemble models. Driven by growing demand, utilities are increasingly exploiting their aging infrastructure towards its operational limits [39], reducing margins and leading to larger size contingency analysis problems [40]. In response, much faster grid analysis functions are required, e.g., including grid contingency events.
Moreover, the lowering of grid inertia through IBR integration (see Section 2.2) is further amplified by the retirement of old synchronous generators. The more frequent use of synchronous generators for regulating purpose bring them more often in operating zones where mechanical or thermal stresses are significant, and thus accelerate the aging of the generators. IBR related switching transients also disrupt existing protection relay schemes and increase grid behavioral uncertainty.
2.5 Cybersecurity threats
DER and IBR technologies are currently being developed and deployed without robust regulatory frameworks mandating essential cybersecurity standards, hence increasing the associated cybersecurity threats. Interconnection and interoperability standards, such as IEEE Std. 1547 requires DERs and IBRs to communicate with clients, vendors, aggregators, and grid operators for monitoring and control. In practice, however, communication often occurs over inadequately secured public and private networks, expanding the cyberattack surface [41]. Consequently, critical vulnerabilities are frequently discovered in commercial DERs and IBRs, such as electric vehicle charging stations [42] and PV inverters [43], increasing the likelihood of major cyberattacks [44]. As the penetration of DERs and IBRs increases, synchronized cyberattacks targeting DERs or IBRs could cause grid reliability issues, from localized instability to system collapse [41]. This risk is exacerbated by reduced grid inertia and contingency reserves due to reliance on zero marginal cost renewables. Additionally, control and management responsibilities for DERs and IBRs are often delegated to aggregators, vendors, and customers. This delegation complicates risk assessment, incident detection, and response – as visibility barriers are created for grid operators – introducing uncertainty and black-box behavior for large sections of their grid. With the lack of real-time data, these barriers create significant additional computational burdens for automated and accelerated data analysis to search for fraudulent injections.
3 Foundation models
Figure 2: Evolution in AI and ML methods, leading to the emergence of FMs
Having identified the growing gaps between computational capabilities and industry needs amidst the energy transition, we now introduce FMs, which is the focus of this article, laying the groundwork for our vision of GridFM.
3.1 The foundation model approach
To understand the FM approach, it helps to review the evolution of AI and ML as depicted in Figure 2. Early hand-crafted symbolic expert systems, from the 1960s onwards, were limited and brittle. The 90s brought big data and general-purpose ML that, however, depend on task-specific hand-crafted features. In 2012, deep learning, enabled by increased compute performance, started to disrupt the AI world [45], but remained task-specific and still required large annotated datasets. Recently, FMs have emerged. They learn from data through self-supervision [46] and have proven to generalize across many applications. Creating an FM is a three-step process, as shown in Figure 3:
  1. 1.
    Pre-training: Developing a general-purpose representation FM through self-supervised pre-training and a suitable architecture, often based on a transformer model with encoder and decoder components. This step involves a reconstruction task where the model reconstructs masked parts of the data.
  2. 2.
    Fine-tuning: The pre-trained FM is customized for various downstream tasks with minimal labeled data by attaching a task-specific decoder head and performing a few training iterations.
  3. 3.
    Inference: The fine-tuned FM is deployed to allow users to request predictions at low computational cost.
Figure 3: FM life-cycle phases i) pre-training, ii) fine-tuning, iii) inference.
FM development is generally computationally expensive, involving pre-training on datasets of tens of terabytes and models with hundreds of millions or even billions of parameters [47]. However, once developed, FMs can be readily adapted to multiple downstream tasks via fine-tuning with minimal labeled data. Thus, the cost of model pre-training can be spread across numerous applications. These economies of scale make it feasible to employ AI for niche applications, for which developing a dedicated traditional AI model might be prohibitively expensive or technically infeasible.
FM training encodes underlying patterns from the supplied data in model parameters. Successful examples of FMs to date include text-based large language models (LLMs) [47] and weather forecast models [48], in which these patterns are linguistic and physics-driven temporal-spatial relationships, respectively. LLMs have been shown to be applied successfully to certain power system problems [14]. By contrast, our focus here is on a multi-modal GridFM able to handle multiple data modalities intrinsic to power grid operations. Modalities of interest include network topologies, sensor data from phasor measurement units (PMUs), SCADA systems, advanced metering infrastructure (AMI), load, generation, weather, and market data.
3.2 Salient advantages
We next review salient features of FMs that hold significant potential for enhancing power grid applications.
Predicting “tokens”
In transformer models, a token is defined as the smallest fundamental unit of data (e.g., a character sequence or image patch). FMs excel at predicting missing tokens based on context, enabling tasks such as sentence completion, programming assistance, and question answering in the text domain [49]. This capability extends beyond text to fields like time-series, audio, or video, where FMs predict next or missing values [50].
Homogenization and scaling
FMs represent a paradigm shift in AI, offering a scalable and standardized approach and making niche applications commercially viable. This AI-as-a-platform approach promises to simplify operations by allowing a few FMs to replace many specialized models. For example, a single geospatial FM for earth observation has been fine-tuned for tasks like flood segmentation, land-use classification, and heat island regression [51].
Accelerating simulations
FMs can emulate physics-based simulations with impressive accuracy and efficiency, potentially combined with physics-informed machine learning [28]. The efficiency gains originate from AI models directly mapping inputs to outputs, unlike traditional physics-based simulations that solve partial differential equations iteratively. Recent AI models [48] emulate numerical weather prediction models [52] with comparable forecasting skills but up to 4–5 orders of magnitude lower computational cost [53] during inference. These game-changing advances improve the spatial and temporal resolution of forecasts and, combined with AI techniques such as diffusion-based models [54], facilitate the creation of large ensembles for uncertainty estimation. Another salient advantage of using AI for emulating simulations lies in the fact that no a-priori assumptions are required but rather the FM can learn the full physics directly from the data or observations. We argue that exactly these characteristics of FMs are crucial for bridging the computational gaps shown in Figure 1.
3.3 Hurdles to implementation
Here, we discuss potential hurdles associated with implementing FMs for the power grid while suggesting solutions.
Need for data
FMs, like other deep learning models, are data hungry, which raises the question of the sufficiency of available data. For example, the Llama3 LLM, with 70 billion parameters, is trained on 15 trillion tokens of text [55]. By way of comparison, the power sector globally generates multiple times that volume daily [56]. Considering the groundbreaking capabilities of large language models in open-ended systems where context and semantics are not consistently described by mathematical equations, there are reasons for optimism about this technology’s potential for the power grid. However, the question remains whether data sets with the right diversity and quality exist or can be assembled so that the full spectrum of scenarios and situations are covered to encode the desired information in model parameters via a suitable training approach.
Data accessibility
Data accessibility is another potential hurdle to overcome. Grid data are heterogeneous and distributed, are owned by multiple entities, and have special security and privacy measures. A potential solution to this barrier includes privacy-preserving federated learning [57] and secure enclave technologies [58]. Such approaches may enable multiple data owners to train models collaboratively without sharing their local data with a central server, thus preserving privacy and complying with data protection regulations. Additional schemes like differential privacy and homomorphic encryption can be integrated into federated learning frameworks [59], ensuring that sensitive information remains protected during training. Secure enclaves, in contrast, create a trusted execution environment for secure processing of sensitive and proprietary data during FM training. Moreover, simulated data generated from real and synthetic systems can complement and provide scenarios under various conditions, such as extreme weather conditions, component failures, and cascading failure events [13]. However, it is important to consider that electrical networks evolve for various reasons, such as network reinforcement, which changes topologies and operating conditions, or new business models (e.g., local energy communities), which change the operating profiles of assets connected to the network, requiring continual learning of the FM.
Trust
Notwithstanding the scenario in which all necessary data, whether synthetic or real, can be cleared and collected for use, there still remains the age-old question of whether the data and, consequently, a data-driven model can be trusted. In general, power system operators prefer interpretable models over black-box models so that the root causes of choices by advanced tools can be better understood, and heuristic solutions can be developed. While the debate on the trustworthiness of AI will continue, there is evidence that AI models can be as trustworthy as first-principle physics-based models, for specific applications [60]. Referring back to the recent breakthroughs of AI emulators for weather [61], it was demonstrated that AI could indeed encode the actual dynamics of certain physical systems and generalize to unseen events in ways that allow it to outperform a physics-based simulation-based model [62]. Moreover, standardized formal verification methods are required for AI. ISO already developed the first family of standards in that direction, e.g., ISO/IEC 24029-2 [63].
The European Union’s recently published AI Act [64] defines two concepts to boost trust in AI-based systems and data sharing: a) testing and experimentation facilities that combine physical and virtual environments to evaluate and certify their latest AI-based software and hardware, and b) the data spaces concept to remove barriers to data sharing, such as interoperability, trust, and privacy, while keeping sovereignty and full control of data.
Malicious use
FMs can be used as a sandbox to accelerate the development of adversarial attacks. For example, FMs’ potential for scalable data imputation, combined with their ability to predict grid operating conditions, could be exploited by adversaries to infer sensitive information about a target power system [65]. Therefore, it is imperative to continuously assess the extent to which misusing such technology can augment malicious actors’ attack capabilities and to adapt cybersecurity defenses and policies in response [66].
4 A concept of foundation model for the power grid
In the previous section, we outlined the salient qualities of FMs by pointing to their established applications in language and weather. Here, we discuss how the challenges in the future power system open up new opportunities for FMs. FMs can reduce computational burdens, accelerating simulations and thus enabling the transition to a sustainable and reliable power system. These models excel in handling complex power grid tasks in near real-time and provide more accurate forecasting or reconstruction of key variables than classical methods. Such problems are well-suited for FMs as they fundamentally involve token prediction or reconstruction tasks. Additionally, an ML agent can be trained to act as an expert in decision-making processes, such as grid expansion, network reconfiguration, and other tasks that often require computationally intensive solutions. Here, we explore the potential implications of developing a GridFM, examining the important components and steps to build such a model.
4.1 Data preprocessing and token design
The successful construction of FMs requires the careful design of tokens to represent data elements effectively. Word and image tokens are essential for successful natural language processing and computer vision training. While these domains have established tokenization methods, strategies must be explored for other domains. For example, graph-based models define a graph vocabulary [67] to capture structural and semantic properties, while time series models segment data into subseries-level patches [68].
Designing tokenization strategies for non-text and non-visual power grid data can be difficult, because the grid involves very heterogeneous, multi-modal datasets, sensor time series, categorical equipment status data, and spatial grid topology data. The PyPSA-Eur dataset [69] is an excellent example of this complexity, where researchers provide a platform to access multi-modal data for the European power system, i.e., network data, electrical demand time series, renewable time series, and geographical potential for solar and wind generations. A promising tokenization strategy for multi-modal data was demonstrated for vision tasks by the massively multi-modal masked modeling architecture (4M) [70]. In this work, good scalability was achieved by unifying the representation space of all modalities, by mapping them into discrete tokens, and performing multi-modal masked modeling on a small randomized subset of tokens. In addition to tokenization, other preprocessing techniques are essential for training models on grid data. For example, normalization methods are critical for training time series models to avoid distribution shift issues [68, 71]. Nevertheless, the alarm data in power grids resembles tokens and natural language elements [72], making FMs applicable for forecasting alarm sequences and filling in missing alarm data.
4.2 Model architecture
Our GridFM model must work with data from at least four modalities: tabular, time series, graph (static and spatiotemporal), and geospatial. Additionally, we expect the power grid dynamics to require a similar number of trainable parameters than weather and climate systems, which are also governed by highly non-linear and complex physical laws. Therefore, an estimate of the number of trainable parameters for GridFMs can be inferred from the FMs developed for these systems, which currently have a few hundred million parameters [53, 73].
We envision two designs for a GridFM that account for the complexity of the grid and its varying transmission and distribution network properties, covering system dynamics from operations to planning. Firstly, it can be designed as a single multi-modal spatiotemporal model. Alternatively, it can consist of multiple FMs integrated into a mixture-of-experts (MoE) framework [74].
The topology of the power grid and its physical properties align well with the message-passing mechanisms used in graph neural networks (GNNs). GNNs are designed for graph-structured data and, through their message-passing mechanisms, learn representations of graph structures called embeddings. These embeddings are learned by iteratively aggregating the features of neighboring nodes. Unlike other deep learning methods that either fail to generalize or require complex data preprocessing, GNNs can generalize to different grid network topologies. Selecting a suitable GNN architecture, however, is challenging. It was argued that the power grid’s unique features necessitate GNNs with depths of up to 27 layers to effectively solve the power flow problem  [75]. More recently, on the other hand, it was demonstrated that more shallow GNNs can achieve very good results for power flow, optimal power flow, and cascading failure analysis on various benchmark transmission grids [13]. Specifically, it was shown that graph attention networks  [76] and transformer-based models perform the best. Attention-based models on graphs assign the importance of a node to its neighbors by learning a score for each node pair, which allows for a more effective understanding of temporal and geographical patterns. Furthermore, attention-based methods such as GraphAny [77] demonstrate high transferability and generalizability to unseen graph topologies for foundation models. Given these insights into the effectiveness of GNNs for power grid analysis, exploring more advanced modeling strategies becomes essential to capture the complexities of the multi-level structure of the power grid. A promising strategy is a hierarchical model architecture, e.g., using GNNs and sublevel time series modeling [68]. Hierarchical GNNs [78] allow for modeling transmission and distribution-level grid topologies. Sublevel time series models can train the FM with various temporal dynamics and resolutions. In addition, hierarchical GNNs can be combined with geometric deep learning for spatiotemporal graphs [79]. The combined use of these sophisticated approaches requires further exploration. It is important to note that the graph representations can go beyond the physical grid topology [80].
Here, we propose a concept of a multi-modal FM for power grids as shown in Figure 4. The model reads structured graphs in which the grid topology is mapped to nodes and edges using GNNs as encoders. Inputs serve as embeddings, including static (tabular and static graphs) and dynamic data (time series, spatiotemporal graphs, and geospatial data). The model also reads the input from variations of the same or different power grids. The input data is partially masked to train an auto-encoder for reconstructing the inputs. The model outputs the reconstructed inputs: measured signals, topology, and simulated data. While using a self-supervised learning approach to reconstruct signals is established, topology reconstruction entails novel techniques. A promising candidate for such a task is the integration of a link prediction approach into the GridFM model. Such an approach allows the model to learn the presence of a masked link using the observed (masked) graph topology [81].
Figure 4: The GridFM concept, which reconstructs power grid-relevant data. Left: Input. Middle: Model Architecture. Right: Output. Grayed-out regions indicate masked tokens. We present examples of inputs to be reconstructed by the foundation model, including time-series signals at the bus, branch, and system levels, tabular data from load flow simulations, and topology reconstructions. Such a model aims to handle various configurations of the same power grid and entirely different power grids.
4.3 Downstream tasks
Once trained, the GridFM can address multiple downstream tasks, with or without additional encoding layers or fine-tuning. We summarize in Figure 5 and discuss the following eight classes of such tasks.
Figure 5: Downstream tasks: Examples of power system problems that a GridFM can address.
Transient and dynamic stability
Here, the GridFM is employed to estimate how a system may respond to disturbances, analyzing signals in both time and frequency domain. The GridFM, when fine-tuned to reconstruct such signals, can help with missing data or predictive modeling. Foundation models with physics-informed learning can efficiently solve the differential algebraic equations governing rotor angle and voltage stability [82]. This can significantly speed up the computation of stability margins and help identify potential stability issues. Moreover, the rapid computation of fault currents using Foundation Models can aid operators in accurately localizing and classifying faults, enabling faster system restoration and minimizing the impact of disturbances on the power system [83].
Load flow
In power system operation and planning, load flow analysis determines the electrical state of the power system, e.g., voltage and phase angles at buses and power flows in branches. These analyses benefit from AI/ML methods, significantly decreasing computational burdens [13]. Traditionally, the power flow and optimal power flow problems are solved with iterative techniques. Due to their non-linear nature and difficulties in converging for large problems, simplifications are often used. Scientific ML techniques have been shown to provide solutions without employing simplified equations, competing with traditional methods in terms of accuracy while being drastically faster [84]. Therefore, with a significant computational speed-up, GridFM has great potential to provide more accurate and scalable solutions for power systems [82].
Forecasting
Forecasting is used in power grids for multiple tasks, including demand prediction, renewable production forecasting, and electricity price prediction. Demand and renewable forecasting are vital in ensuring power generation-demand balance, which is becoming increasingly important in a grid with high renewable penetration. The self-supervised learning paradigm used in FMs and their predictive ability make them intrinsically suitable for forecasting [85]. Moreover, FMs can integrate exogenous variables, such as numerical weather predictions, which also carry associated forecast uncertainties.
Control operations
Here, the concerns include frequency and voltage control, network reconfiguration management, protection activation, and load shedding. FMs will help optimize and control the power system through sequential decision-making agents [86]. A GridFM can enhance reinforcement learning (RL) agents by introducing prior knowledge of the environment before and during training, improving learning efficiency, and reducing risks.
Electricity markets
Electricity markets will benefit from the computational efficiency and predictive ability of a GridFM. The different markets, including futures, balancing, day-ahead, and intraday markets, can be optimized using the base GridFM or fine-tuned versions. Furthermore, the model will identify the security-based re-dispatch after the day-ahead market clearing in a timely manner [87]. It will also help to pinpoint the location, amount, and type of ancillary services from distributed sources for congestion management in the future power grid with high shares of renewables. Finally, a GridFM will provide prompt solutions for economic dispatch and unit commitment problems, traditionally solved as linear and mixed-integer optimization problems [88].
System security
ML methods have been used to identify faulty components [12] and to speed up simulations like $N-1$, $N-K$ contingency, and cascading failure analyses [89], all critical tasks for operation and planning. Such models are faster and better leverage past knowledge for security predictions than traditional tools. Nevertheless, these analyses require additional encoding to incorporate topology with bus or branch-level features to inform about the final state of the entire grid. In this context, a GridFM tuned for system security will explore a broader and more complete set of scenarios, and thus both aid in designing a reliable power system and provide a system security tool for power system operations planning.
Expansion planning
Rapid electrification of the industrial and automotive sectors and climate change’s impact on generation profiles make expansion planning a crucial task at distribution and transmission levels. The generative capabilities of a GridFM to create future scenarios for load and generation (e.g., using public projections of electric vehicle sales and national renewable energy plans) will enhance decision-making. By integrating scenario generation with graph generation and deep reinforcement learning techniques, optimized and upgraded power systems can be designed to capture a broader spectrum of future extreme scenarios [90]. Additionally, by exploiting cross-country information (e.g., the impact of interconnection flows), a GridFM will improve expansion planning strategies to prepare for these effects and enable the energy transition in the following decades.
Cybersecurity
Cybersecurity challenges could be tackled by leveraging FMs to speed up power system simulations. These FMs facilitate risk assessment and incident response in transmission and distribution systems by identifying and prioritizing threat scenarios [91] and designing optimal responses to maintain grid reliability [92]. They can generate synthetic cybersecurity data by learning complex power system dynamics to overcome data scarcity, improving intrusion detection, and enabling dynamic stress testing against intelligent adversaries [93]. A cyber-physical FM can combine SCADA and AMI data with cybersecurity information to enhance cyber threat mitigation, detection, and response.
5 Conclusion
The rapid emergence of FMs is reshaping the AI landscape. While such models are well established for text and image analysis, here we have introduced the idea of a GridFM: an FM for the power grid, able to learn from time-series, geospatial, and other data to improve power grid planning, operations, and control. We first motivated the need for a GridFM by pointing to the growing gaps between computational capabilities and the industry’s need to cope with the increasing uncertainty and complexity of the power grid, largely due to the energy transition. We then reviewed the three core advantages of the FM approach for power systems: excellent missing token prediction skills (signal reconstruction); broad adaptability and scalability; and simulation acceleration. In another domain characterized by apparently intractable complexity, namely weather forecasting, we have seen AI weather FMs approximate the fundamental equations governing weather to produce much more computationally efficient models (more than 4 orders of magnitude speed-up).
In a similar manner, we propose to develop a GridFM that learns the behaviors and dynamics from measurements and simulations for different network configurations. The superior computational performance of such a GridFM has the potential to accelerate grid analyses by several orders of magnitudes. This acceleration will, in turn, enable modeling of the increasingly complex (e.g., bidirectional power flow, combined transmission, and distribution) and uncertain grid (e.g., by modeling different scenarios going beyond the traditional N-1 analysis). The promise of such a GridFM is that once pre-trained on multiple grids and various networks, it can be readily fine-tuned to specific grid topologies (distribution and transmission). In summary, we argue that research in the proposed FM for the power grid is imperative to facilitate the energy transition. We envision the power system and AI communities collaborating to take on this “moonshot” and thus revolutionize power grid analysis, leveraging AI advancements to meet common goals and needs.
Acknowledgement
The work was supported by the respective organizations of the authors. The authors acknowledge the invaluable discussions and interactions in the working group for Foundation Models for the Power Grid, including Roya Ahmadi, Rich Argentieri, Gianni Vittorio Armani, Venkat Balachandran, Ranjini Bangalore, Venkat Banunarayanan, Edmon Begoli, Larry Bekkedahl, Keith Benes, Michelle Berti, Budhu Bhaduri, Roberto Bianchessi, Gilbert Bindewald, Louise Blais, David Bromberg, Matthias Bucher, Eren Cam, Supriya Chinthavali, Seong Choi, Julien Chosson, Joseph Collum, Magnus Cormack, Peter De Bock, Fei Ding, Kathy Duviella, Darlain Edeme, Yamshid Farhat, Mark Fenton, Marcus Freitag, Alirez Ghassemian, Tassos Golnas, Gavin Goodland, Luca Grella, John Grosh, Benois Grossin, Andy Hackett, Hagen Haentsch, Zeryai Hagos, Grahame Hay, Frank Hellmann, Andy Henton, Raffael Hilber, Dan Hilton, Gabriela Hug, Alex Hurley, Sorana Ionescu, Kevin Johnston, Gareth Jones, Adrian Kelly, Bernd Klockl, Teddy Kott, Benjamin Kroposki, Teja Kuruganti, Larry Light, Dalton Lunga, Matt Magill, Adrian Maldonado, Antoine Maret, Pete Marsh, Russel Mason, James McClean, John Messer, Nikki Militello, Jordan Murkin, Xabi Muruaga, Indu Nambiar, Christian Nauck, Caleb Northrop, Gabby Nyirjesy , Marco Padula, Patrick Panciatici, Raul Alfaro Pelico, Mark Petri, Pierre Pinson, Abhishek Pontis, Josh Power, Alan Pressman, Feng Qiu, Mark Rafferty, Pippa Robertson, Nicoletta Rocca, Caroline Rose-Newport, Lyndon Ruff, Bryan Sacks, Kelly Sanders, Giovanni Sansavini, Michael Sheppard, Robert Shorten, Vikas Singhvi, Stanislav Sobolevsky, Pratik Sonthalia, Brian Spears, Declan Stock, Olle Sundstroem, Matthew Tarduogno, Roger Thompson, Anna Carolina Tortora, Alex Vivas, Lucy Vu, Ben Weiner, Casey Werth, Daniel Wetzel, Gary White, Xin Wu, Steven Wysmuller, Yunchi Yang, Thomas Zadlo, Marzia Zafar and many others. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.
Author contributions
Conceptualization, H.F.H., L.S.A.M. and J.W.; Investigation, H.F.H., T.B., B.G., L.S.A.M., A.P., A.V., J.W., J.B., A.B.M., S.C., I.F., B.H., R.J., K.K., V.M., F.M., M.D.M., O.R., H.S., L.X., E.S.Y., and A.Z.; Writing – Original Draft, H.F.H., T.B., B.G., L.S.A.M., A.P., A.V., J.W., J.B., A.B.M., S.C., I.F., B.H., R.J., K.K., V.M., F.M., M.D.M., O.R., H.S., L.X., E.S.Y., and A.Z.; Writing – Review & Editing, A.J.B., R.J.B., B.P.B., J.S. and S.S.; Project Administration, H.F.H.
Declaration of Interests
The authors declare no competing interests. H.F.H is also affiliated with Yamagata University and University of Illinois Urbana-Champaign. S.S. is also affiliated with Masaryk University.

### v2

@@
-A Perspective on Foundation Models for the Electric Power Grid
-Abstract
-Foundation models (FMs) currently dominate news headlines.
-They employ advanced deep learning architectures to extract structural information autonomously from vast datasets through self-supervision.
-The resulting rich representations of complex systems and dynamics can be applied to many downstream applications.
-Therefore, FMs can find uses in electric power grids, challenged by the energy transition and climate change.
-In this paper, we call for the development of, and state why we believe in, the potential of FMs for electric grids.
-We highlight their strengths and weaknesses amidst the challenges of a changing grid.
-We argue that an FM learning from diverse grid data and topologies could unlock transformative capabilities, pioneering a new approach in leveraging AI to redefine how we manage complexity and uncertainty in the electric grid.
-Finally, we discuss a power grid FM concept, namely GridFM, based on graph neural networks and show how different downstream tasks benefit.
-keywords:
-Foundation Models , Data-Driven Power Grid Modeling , Energy Transition , AI-based Power Flow Simulation
+Foundation Models for the Electric Power Grid
+Foundation Models for the Electric Power Grid
+Hendrik F.
+Hamann§* Blazhe Gjorgiev* Thomas Brunschwiler* Leonardo S.
+A.
+Martins* Alban Puech* Anna Varbella* Jonas Weiss* Juan Bernabe-Moreno Alexandre Blondin Massé Seong Choi Ian Foster Bri-Mathias Hodge†, Rishabh Jain Kibaek Kim Vincent Mai François Mirallès Martin De Montigny Octavio Ramos-Leaños Hussein Suprême Le Xie El-Nasser S.
+Youssef Arnaud Zinflou Alexander Belyi Ricardo J.
+Bessa Bishnu Prasad Bhattarai Johannes Schmude Stanislav Sobolevsky
 […]
@@
 [bra]organization=Masaryk University,addressline=Kotlarska, 2, city=Brno, postcode=62000, country=Czech Republic
-footnotetext: § Lead contact: hehamann.hh@gmail.com or hendrikh@us.ibm.comfootnotetext: * These authors contributed equally to this work as major contributors.footnotetext: † These authors contributed equally to this work as significant contributors.footnotetext: ‡ These authors contributed equally to this work.
+§ Lead contact: hehamann.hh@gmail.com
+* These authors contributed equally to this work as main contributors.
+† These authors contributed equally to this work as significant contributors.
+‡ These authors contributed equally to this work.
+Keywords
+Foundation Models, Data-Driven Power Grid Modeling, Energy Transition, AI-based Power Flow Simulation
+Summary
+Foundation Models (FMs) currently dominate news headlines.
+They employ advanced deep learning architectures to extract structural information autonomously from vast data-sets through self-supervision.
+The resulting rich representations of complex systems and dynamics can be applied to many downstream applications.
+Therefore, advances in FMs can find uses in electric power grids, challenged by the energy transition and climate change.
+This paper calls for the development of FMs for electric grids.
+We highlight their strengths and weaknesses amidst the challenges of a changing grid.
+It is argued that FMs learning from diverse grid data and topologies, that we call GridFMs, could unlock transformative capabilities, pioneering a new approach in leveraging AI to redefine how we manage complexity and uncertainty in the electric grid.
+Finally, we discuss a practical implementation pathway and road map of a GridFM–v0, a first GridFM for power flow applications based on graph neural networks, and explore how various downstream use cases will benefit from this model and future GridFMs.
 1 Introduction
@@
 This is creating a widening gap between existing computational capabilities and the evolving needs of the electric power industry.
-AC PF denotes alternating current power flow; DC PF denotes direct current power flow.
 In this context, it is clear that the operation, control, and planning of power systems will soon be pushed to their limits.
@@
 Machine Learning (ML) and Artificial Intelligence (AI) methods have shown promise for such purposes in a wide spectrum of industries, with significant breakthroughs in computer vision [8], natural language processing [9], and intelligent control [10].
-Such approaches rely on statistical algorithms to learn from massive data, often enabling them to execute tasks (e.g., regression, classification, detection, segmentation, forecasting, generation) without human supervision.
-Yet, despite significant progress in the application of AI/ML methods to power systems [11, 12, 13], wider adoption has been impeded, in particular, by a lack of readily available training data and the limited transferability of developed models to adjacent applications [14].
+Such approaches rely on statistical algorithms to learn from massive data, often enabling them to execute tasks (e.g., regression, classification, detection, segmentation, forecasting, and generation) without human supervision.
+Yet, despite significant progress in the application of AI/ML methods to power systems [11, 12, 13], wider adoption has been impeded, in particular, by a lack of readily available training data and the limited transferability of developed models to adjacent applications and system configurations [14].
 This is where, we argue, Foundation Models (FM) have an important role to play.
-FMs are advanced AI models developed through self-supervised learning, most often based on transformer architectures, that generalize across various tasks after initial training on large datasets, enabling efficient adaptation to specific applications with minimal labeled data.
+FMs are advanced AI models developed through self-supervised learning, most often based on transformer architectures, that generalize across various tasks after initial training on large data-sets, enabling efficient adaptation to specific applications with minimal labeled data.
 Text-based FMs like GPT-4 [15], encoding deep relationships in text, have enabled a wide range of new applications.
-Similarly, we anticipate that a multi-modal FM for power grids, trained on the diverse data that characterizes the power grid, will enable a broad set of new use cases.
-To enable this, we advocate in this article for developing multi-modal FMs for power grids: what we refer to as GridFMs.
+Similarly, we anticipate that FMs for power grids, trained on the diverse data that characterizes the power grid, will enable a broad set of new use cases.
+In this article, we advocate for developing FMs for power grids: what we refer to as GridFMs.
 We explain why we believe in their potential, highlighting their anticipated strengths and weaknesses in light of the challenges posed by the evolving power grid.
-We develop a concrete picture of what we envision a GridFM to be and how we expect it to pioneer a new approach in leveraging AI to manage complexity and uncertainty in the power grid.
+We introduce a practical road map for implementing GridFM–v0, an initial version of a GridFM for power flow applications, and discuss our vision for future GridFMs, pioneering a new approach in leveraging AI to manage complexity and uncertainty in the power grid.
 In the following, we first highlight, in Section 2, the looming power grid challenges associated with the energy transition, which are further aggravated by aging infrastructure, swelling cybersecurity, and climate change concerns.
 We show how, by increasing the complexity and uncertainty of the power grid, these drivers are creating a significant gap in computational capabilities, which we propose to overcome with FMs.
-We then review in Section 3 the strengths and weaknesses of FMs and propose in Section 4 an implementation of a GridFM that we argue will enable exciting, currently unattainable use cases as downstream tasks.
+We then review in Section 3 the strengths and weaknesses of FMs and propose in Section 4 the near-term and long-term development phases of GridFMs that we argue will overcome current limitations and enable scenarios, which are currently unattainable.
 2 Looming challenges
@@
 While a single coal plant can generate hundreds of megawatts predictably and reliably, large numbers of weather-dependent renewable energy resources like wind and solar are required to produce the same power output.
-Current hourly and daily forecast methods still show up to 3% and up to 10% forecast error, respectively [16].
+Current methods to forecast 1-day or 1-hour ahead still show up to 3% to 10% forecast error, respectively [16], depending on the energy source.
 Additionally, Distributed Energy Resources (DERs) like battery storage and electric vehicles are operated by independent power management systems and are driven by user patterns with intrinsic uncertainties.
@@
 The numbers of controllable and dispatchable devices, as well as balancing nodes are expected to grow exponentially [17] with additional thousands to millions to be added in the near future.
-Managing the associated variability and uncertainty will require advanced stochastic optimization, scenario reduction, and sampling-based approaches, which are computationally more intensive than currently used deterministic methods, especially for real-time demand-supply balancing [18].
-Furthermore, as the net load of DER-rich distribution branches seen by transmission operators becomes more variable and less predictable, effective grid management and planning will become impossible without effective co-simulation frameworks that integrate transmission and distribution analysis.
+Managing the associated variability and uncertainty will require advanced stochastic optimization, scenario reduction, and sampling-based approaches, which are computationally more intensive than currently used deterministic methods, especially for real-time demand-supply balancing.Furthermore, as the net load of DER-rich distribution branches seen by transmission operators becomes more variable and less predictable, effective grid management and planning will become impossible without effective co-simulation frameworks that integrate transmission and distribution analysis.
 Yet the computational cost of such simulations, at least with existing technologies, is extremely high.
@@
 With more inverters connected, grid inertia – which stabilizes phase, voltage, and frequency during transient events and which is mostly tied to large synchronous generators – decreases [18], unless more grid-forming converters are introduced.
-As inverters are driven by software, they can quickly adjust power or disconnect, causing additional issues like sub-synchronous resonances and even trip large generators [20].
-Despite the increasing adoption of standards [21], inverters remain hard to predict and model, unlike physics-driven synchronous generators.
-Quasi-static phasor simulations are used to assess optimal power flow.
-For grid stability assessment, however, computationally expensive electro-magnetic transient (EMT) simulations [22] are required.
-They run mostly on traditional central processing units (CPUs), which are flexible for many algorithms but inferior in compute performance to recent graphics processing units (GPUs), which are just starting to get adopted in the industry, as, e.g., in [23].
-Also, CPU systems do not always scale well for large EMT simulations [24].
-Grid operations, however, would require EMT simulations to run in real-time and for large sections of the grid, including distribution grid sections.
-To further account for inherent network uncertainties and those introduced by inverter-based resources (IBRs), stochastic ensembles of the problems would need to be simulated.
-Altogether, such an approach is computationally expensive and, with current approaches, impractical [25].
-Thus, the community mostly relies on offline simulators like PS CAD [26] or ParaEMT [27] while also exploring physics-informed neural networks to reduce EMT simulation time [28].
+Because inverters are controlled by software, they can quickly adjust power or disconnect, causing additional issues like sub-synchronous resonances and even trip large generators [19].
+[…]
+To further account for inherent network uncertainties and those introduced by Inverter-Based and Distributed Energy Resources (IBRs, DERs), stochastic ensembles of the problems must be simulated.
+Altogether, such an approach is computationally very expensive and, with current approaches, impractical [25].
+Thus, the community mostly relies on offline simulators like PS CAD [26] or ParaEMT [23] while starting exploring physics-informed neural networks (PINNs) to reduce EMT simulation time [27].
 2.3 Changes to demand and weather patterns
@@
 To improve model accuracy, additional information that captures demographic changes and changes in the distribution of wealth, which lead to different societal and individual behaviors (e.g., heating, cooling, electromobility), as well as climate change-induced environmental shifts must be integrated.
-Therefore, the capacity to combine heterogeneous data sources (including information in textual format) is a fundamental requirement to improve predictability.
+Therefore, the capacity to combine heterogeneous data sources, including information in textual format, is a fundamental requirement to improve predictability.
 2.4 Aging infrastructure
@@
 The more frequent use of synchronous generators for regulating purpose bring them more often in operating zones where mechanical or thermal stresses are significant, and thus accelerate the aging of the generators.
-IBR related switching transients also disrupt existing protection relay schemes and increase grid behavioral uncertainty.
+IBR related switching transients also disrupt existing protection relay schemes [40] and increase grid behavioral uncertainty.
 2.5 Cybersecurity threats
@@
 Figure 2: Evolution in AI and ML methods, leading to the emergence of FMs
-Having identified the growing gaps between computational capabilities and industry needs amidst the energy transition, we now introduce FMs, which is the focus of this article, laying the groundwork for our vision of GridFM.
+Having identified the growing gaps between computational capabilities and industry needs amidst the energy transition, we now introduce FMs, laying the groundwork for our vision of GridFMs.
 3.1 The foundation model approach
-To understand the FM approach, it helps to review the evolution of AI and ML as depicted in Figure 2.
+To understand the FM approach, it is helpful to review the evolution of AI and ML as depicted in Figure 2.
 Early hand-crafted symbolic expert systems, from the 1960s onwards, were limited and brittle.
-The 90s brought big data and general-purpose ML that, however, depend on task-specific hand-crafted features.
-In 2012, deep learning, enabled by increased compute performance, started to disrupt the AI world [45], but remained task-specific and still required large annotated datasets.
+The 90s brought big data and general-purpose ML that, however, still often depend on task-specific hand-crafted features.
+In 2012, deep learning, enabled by increased compute performance, started to disrupt the AI world [45], but was more often than not used to train models on large annotated data-sets in a supervised way, making them task-specific.
 Recently, FMs have emerged.
-They learn from data through self-supervision [46] and have proven to generalize across many applications.
-Creating an FM is a three-step process, as shown in Figure 3:
+They learn from data through self-supervision and large content-rich data-sets [46] and have proven to generalize across many applications.
+FMs draw their potential from being able to train on very large data-sets, which have not been annotated, thus are more easily available than labeled data-sets, as required for supervised learning.
+The FM approach is a three-step process, as shown in Figure 3:
   1. 1.
-    Pre-training: Developing a general-purpose representation FM through self-supervised pre-training and a suitable architecture, often based on a transformer model with encoder and decoder components.
+    Pre-training: A general-purpose FM is developed through self-supervised pre-training and a suitable architecture, often based on a transformer model with encoder and decoder components.
 This step involves a reconstruction task where the model reconstructs masked parts of the data.
   2. 2.
-    Fine-tuning: The pre-trained FM is customized for various downstream tasks with minimal labeled data by attaching a task-specific decoder head and performing a few training iterations.
+    Fine-tuning: The pre-trained FM is customized for various downstream tasks and applications with minimal labeled data by attaching a task-specific decoder head and performing a few training iterations.
   3. 3.
@@
 Figure 3: FM life-cycle phases i) pre-training, ii) fine-tuning, iii) inference.
-FM development is generally computationally expensive, involving pre-training on datasets of tens of terabytes and models with hundreds of millions or even billions of parameters [47].
-However, once developed, FMs can be readily adapted to multiple downstream tasks via fine-tuning with minimal labeled data.
+The development of FMs involves extensive pre-training on data-sets of tens of terabytes and models with hundreds of millions or even billions of parameters [47].
+However, in contrast to what one might assume, even bespoke AI models can be very complex and expensive to train.
+For example, the non-foundation and application-specific weather AI models, Graphcast (for forecasting) or CorrDiff (for downscaling) were trained on 32 TPU devices [48] and on 128 A100 GPUs [49], respectively.
+The key advantage of FMs is that once developed, they can be readily adapted to multiple downstream applications via fine-tuning with minimal labeled data.
 Thus, the cost of model pre-training can be spread across numerous applications.
 These economies of scale make it feasible to employ AI for niche applications, for which developing a dedicated traditional AI model might be prohibitively expensive or technically infeasible.
-FM training encodes underlying patterns from the supplied data in model parameters.
-Successful examples of FMs to date include text-based large language models (LLMs) [47] and weather forecast models [48], in which these patterns are linguistic and physics-driven temporal-spatial relationships, respectively.
+FM pre-training encodes underlying patterns from the supplied data in model parameters.
+Successful examples of FMs to date include text-based Large Language Models (LLMs) [47] but also weather forecast models [50], in which these patterns are linguistic and physics-driven temporal-spatial relationships, respectively.
 LLMs have been shown to be applied successfully to certain power system problems [14].
-By contrast, our focus here is on a multi-modal GridFM able to handle multiple data modalities intrinsic to power grid operations.
-Modalities of interest include network topologies, sensor data from phasor measurement units (PMUs), SCADA systems, advanced metering infrastructure (AMI), load, generation, weather, and market data.
+By contrast, our focus here is on GridFMs that go beyond text, that will eventually be able to handle multiple data modalities intrinsic to power grid operations.
+Modalities of interest include network topologies, sensor data from Phasor Measurement Units (PMUs), alarms, SCADA systems, Advanced Metering Infrastructure (AMI), load, generation, weather, and market data.
 3.2 Salient advantages
-We next review salient features of FMs that hold significant potential for enhancing power grid applications.
+We next review salient features of AI in general and FMs in particular that hold significant potential for enhancing power grid applications.
 Predicting “tokens”
 In transformer models, a token is defined as the smallest fundamental unit of data (e.g., a character sequence or image patch).
-FMs excel at predicting missing tokens based on context, enabling tasks such as sentence completion, programming assistance, and question answering in the text domain [49].
+FMs have been shown to excel at predicting missing tokens based on context, enabling tasks such as sentence completion, programming assistance, and question answering in the text domain [51].
 This capability extends beyond text to fields like time-series, audio, or video, where FMs predict next or missing values [52].
@@
 Accelerating simulations
-FMs can emulate physics-based simulations with impressive accuracy and efficiency, potentially combined with physics-informed machine learning [28].
+AI models in general, including FMs can emulate physics-based simulations with impressive accuracy and efficiency, potentially combined with physics-informed machine learning [27].
 The efficiency gains originate from AI models directly mapping inputs to outputs, unlike traditional physics-based simulations that solve partial differential equations iteratively.
@@
 These game-changing advances improve the spatial and temporal resolution of forecasts and, combined with AI techniques such as diffusion-based models [57], facilitate the creation of large ensembles for uncertainty estimation.
-Another salient advantage of using AI for emulating simulations lies in the fact that no a-priori assumptions are required but rather the FM can learn the full physics directly from the data or observations.
-We argue that exactly these characteristics of FMs are crucial for bridging the computational gaps shown in Figure 1.
+Another salient advantage of using AI for emulating simulations lies in the fact that no a-priori assumptions are required, but rather the AI model can learn the full physics directly from the data or observations.
+We argue that exactly these characteristics of AI in general and FMs in particular are crucial for bridging the computational gaps shown in Figure 1.
 3.3 Hurdles to implementation
@@
 Need for data
-FMs, like other deep learning models, are data hungry, which raises the question of the sufficiency of available data.
+FMs, like other deep learning models, are data hungry.
 For example, the Llama3 LLM, with 70 billion parameters, is trained on 15 trillion tokens of text [58].
 By way of comparison, the power sector globally generates multiple times that volume daily [59].
-Considering the groundbreaking capabilities of large language models in open-ended systems where context and semantics are not consistently described by mathematical equations, there are reasons for optimism about this technology’s potential for the power grid.
-However, the question remains whether data sets with the right diversity and quality exist or can be assembled so that the full spectrum of scenarios and situations are covered to encode the desired information in model parameters via a suitable training approach.
+Considering the groundbreaking capabilities of large language models in open-ended systems where context and semantics are not consistently described by mathematical equations, we are confident about this technology’s potential for the power grid.
+Diverse and high-quality data-sets required for training grid FMs will, though, still have to be assembled to cover the full spectrum of scenarios and situations.
+Addressing transmission grid problems first, where data is available in a more standardized form, e.g., PMU or power flow data will help demonstrate the value of FMs for the grid.
+This will create incentives for data at the distribution level, which currently is much more heterogeneous and of less quality [60], to be collected, curated, and made available in suitable formats for use with AI and FMs at a larger scale.
 Data accessibility
-Data accessibility is another potential hurdle to overcome.
+Data accessibility is another hurdle to overcome.
 Grid data are heterogeneous and distributed, are owned by multiple entities, and have special security and privacy measures.
@@
 Secure enclaves, in contrast, create a trusted execution environment for secure processing of sensitive and proprietary data during FM training.
-Moreover, simulated data generated from real and synthetic systems can complement and provide scenarios under various conditions, such as extreme weather conditions, component failures, and cascading failure events [13].
+Moreover, simulated data from real and synthetic systems can complement and provide scenarios under various conditions, such as extreme weather conditions, component failures, and cascading failure events [13].
+As discussed further in Section 4, GridFMs might be developed by synthetic data, which are then updated and verified with proprietary data behind the user’s firewall.
 However, it is important to consider that electrical networks evolve for various reasons, such as network reinforcement, which changes topologies and operating conditions, or new business models (e.g., local energy communities), which change the operating profiles of assets connected to the network, requiring continual learning of the FM.
@@
 In general, power system operators prefer interpretable models over black-box models so that the root causes of choices by advanced tools can be better understood, and heuristic solutions can be developed.
-While the debate on the trustworthiness of AI will continue, there is evidence that AI models can be as trustworthy as first-principle physics-based models, for specific applications [60].
-Referring back to the recent breakthroughs of AI emulators for weather [61], it was demonstrated that AI could indeed encode the actual dynamics of certain physical systems and generalize to unseen events in ways that allow it to outperform a physics-based simulation-based model [62].
-Moreover, standardized formal verification methods are required for AI.
-ISO already developed the first family of standards in that direction, e.g., ISO/IEC 24029-2 [63].
+While the debate on the trustworthiness of AI will continue, there is evidence that AI models can be as performant as first-principle physics-based models, at least for specific applications [64].
+Performance alone however does not seem to be necessarily a sufficient condition for trust because even conventional solvers can have stability and convergence challenges [65], which we have found ways to deal with through decades of usage.
+Referring back to the recent breakthroughs of AI emulators for weather, it was demonstrated that AI could indeed encode the actual dynamics of certain physical systems and generalize to unseen events in ways that allow it to have comparable performances to a physics-based, simulation-based model [66].
+In some cases, e.g.
+for an AI power flow emulator, which is discussed in Section 4, the AI results can be immediately verified by using the physics equations.
+In other cases, where such ad-hoc verification is not possible, a consistent way is required to evaluate the model for out-of-sample scenarios.
+In the longer-term a set of standards must be developed as to what performances must be achieved for a given out-of sample scenario.
+Towards that end, ISO has already developed the first family of standards, e.g., ISO/IEC 24029-2 [67].
 The European Union’s recently published AI Act [68] defines two concepts to boost trust in AI-based systems and data sharing: a) testing and experimentation facilities that combine physical and virtual environments to evaluate and certify their latest AI-based software and hardware, and b) the data spaces concept to remove barriers to data sharing, such as interoperability, trust, and privacy, while keeping sovereignty and full control of data.
@@
 Therefore, it is imperative to continuously assess the extent to which misusing such technology can augment malicious actors’ attack capabilities and to adapt cybersecurity defenses and policies in response [70].
-4 A concept of foundation model for the power grid
-In the previous section, we outlined the salient qualities of FMs by pointing to their established applications in language and weather.
-Here, we discuss how the challenges in the future power system open up new opportunities for FMs.
-FMs can reduce computational burdens, accelerating simulations and thus enabling the transition to a sustainable and reliable power system.
-These models excel in handling complex power grid tasks in near real-time and provide more accurate forecasting or reconstruction of key variables than classical methods.
-Such problems are well-suited for FMs as they fundamentally involve token prediction or reconstruction tasks.
-Additionally, an ML agent can be trained to act as an expert in decision-making processes, such as grid expansion, network reconfiguration, and other tasks that often require computationally intensive solutions.
-Here, we explore the potential implications of developing a GridFM, examining the important components and steps to build such a model.
-4.1 Data preprocessing and token design
-The successful construction of FMs requires the careful design of tokens to represent data elements effectively.
-Word and image tokens are essential for successful natural language processing and computer vision training.
-While these domains have established tokenization methods, strategies must be explored for other domains.
-For example, graph-based models define a graph vocabulary [67] to capture structural and semantic properties, while time series models segment data into subseries-level patches [68].
-Designing tokenization strategies for non-text and non-visual power grid data can be difficult, because the grid involves very heterogeneous, multi-modal datasets, sensor time series, categorical equipment status data, and spatial grid topology data.
-The PyPSA-Eur dataset [69] is an excellent example of this complexity, where researchers provide a platform to access multi-modal data for the European power system, i.e., network data, electrical demand time series, renewable time series, and geographical potential for solar and wind generations.
-A promising tokenization strategy for multi-modal data was demonstrated for vision tasks by the massively multi-modal masked modeling architecture (4M) [70].
-In this work, good scalability was achieved by unifying the representation space of all modalities, by mapping them into discrete tokens, and performing multi-modal masked modeling on a small randomized subset of tokens.
-In addition to tokenization, other preprocessing techniques are essential for training models on grid data.
-For example, normalization methods are critical for training time series models to avoid distribution shift issues [68, 71].
-Nevertheless, the alarm data in power grids resembles tokens and natural language elements [72], making FMs applicable for forecasting alarm sequences and filling in missing alarm data.
-4.2 Model architecture
-Our GridFM model must work with data from at least four modalities: tabular, time series, graph (static and spatiotemporal), and geospatial.
-Additionally, we expect the power grid dynamics to require a similar number of trainable parameters than weather and climate systems, which are also governed by highly non-linear and complex physical laws.
-Therefore, an estimate of the number of trainable parameters for GridFMs can be inferred from the FMs developed for these systems, which currently have a few hundred million parameters [53, 73].
-We envision two designs for a GridFM that account for the complexity of the grid and its varying transmission and distribution network properties, covering system dynamics from operations to planning.
-Firstly, it can be designed as a single multi-modal spatiotemporal model.
-Alternatively, it can consist of multiple FMs integrated into a mixture-of-experts (MoE) framework [74].
-The topology of the power grid and its physical properties align well with the message-passing mechanisms used in graph neural networks (GNNs).
-GNNs are designed for graph-structured data and, through their message-passing mechanisms, learn representations of graph structures called embeddings.
-These embeddings are learned by iteratively aggregating the features of neighboring nodes.
-Unlike other deep learning methods that either fail to generalize or require complex data preprocessing, GNNs can generalize to different grid network topologies.
-Selecting a suitable GNN architecture, however, is challenging.
-It was argued that the power grid’s unique features necessitate GNNs with depths of up to 27 layers to effectively solve the power flow problem  [75].
-More recently, on the other hand, it was demonstrated that more shallow GNNs can achieve very good results for power flow, optimal power flow, and cascading failure analysis on various benchmark transmission grids [13].
-Specifically, it was shown that graph attention networks  [76] and transformer-based models perform the best.
-Attention-based models on graphs assign the importance of a node to its neighbors by learning a score for each node pair, which allows for a more effective understanding of temporal and geographical patterns.
-Furthermore, attention-based methods such as GraphAny [77] demonstrate high transferability and generalizability to unseen graph topologies for foundation models.
-Given these insights into the effectiveness of GNNs for power grid analysis, exploring more advanced modeling strategies becomes essential to capture the complexities of the multi-level structure of the power grid.
-A promising strategy is a hierarchical model architecture, e.g., using GNNs and sublevel time series modeling [68].
-Hierarchical GNNs [78] allow for modeling transmission and distribution-level grid topologies.
-Sublevel time series models can train the FM with various temporal dynamics and resolutions.
-In addition, hierarchical GNNs can be combined with geometric deep learning for spatiotemporal graphs [79].
-The combined use of these sophisticated approaches requires further exploration.
-It is important to note that the graph representations can go beyond the physical grid topology [80].
-Here, we propose a concept of a multi-modal FM for power grids as shown in Figure 4.
-The model reads structured graphs in which the grid topology is mapped to nodes and edges using GNNs as encoders.
-Inputs serve as embeddings, including static (tabular and static graphs) and dynamic data (time series, spatiotemporal graphs, and geospatial data).
-The model also reads the input from variations of the same or different power grids.
-The input data is partially masked to train an auto-encoder for reconstructing the inputs.
-The model outputs the reconstructed inputs: measured signals, topology, and simulated data.
-While using a self-supervised learning approach to reconstruct signals is established, topology reconstruction entails novel techniques.
-A promising candidate for such a task is the integration of a link prediction approach into the GridFM model.
-Such an approach allows the model to learn the presence of a masked link using the observed (masked) graph topology [81].
-Figure 4: The GridFM concept, which reconstructs power grid-relevant data.
-Left: Input.
-Middle: Model Architecture.
-Right: Output.
-Grayed-out regions indicate masked tokens.
-We present examples of inputs to be reconstructed by the foundation model, including time-series signals at the bus, branch, and system levels, tabular data from load flow simulations, and topology reconstructions.
-Such a model aims to handle various configurations of the same power grid and entirely different power grids.
-4.3 Downstream tasks
-Once trained, the GridFM can address multiple downstream tasks, with or without additional encoding layers or fine-tuning.
-We summarize in Figure 5 and discuss the following eight classes of such tasks.
-Figure 5: Downstream tasks: Examples of power system problems that a GridFM can address.
-Transient and dynamic stability
-Here, the GridFM is employed to estimate how a system may respond to disturbances, analyzing signals in both time and frequency domain.
-The GridFM, when fine-tuned to reconstruct such signals, can help with missing data or predictive modeling.
-Foundation models with physics-informed learning can efficiently solve the differential algebraic equations governing rotor angle and voltage stability [82].
-This can significantly speed up the computation of stability margins and help identify potential stability issues.
-Moreover, the rapid computation of fault currents using Foundation Models can aid operators in accurately localizing and classifying faults, enabling faster system restoration and minimizing the impact of disturbances on the power system [83].
-Load flow
-In power system operation and planning, load flow analysis determines the electrical state of the power system, e.g., voltage and phase angles at buses and power flows in branches.
-These analyses benefit from AI/ML methods, significantly decreasing computational burdens [13].
-Traditionally, the power flow and optimal power flow problems are solved with iterative techniques.
-Due to their non-linear nature and difficulties in converging for large problems, simplifications are often used.
-Scientific ML techniques have been shown to provide solutions without employing simplified equations, competing with traditional methods in terms of accuracy while being drastically faster [84].
-Therefore, with a significant computational speed-up, GridFM has great potential to provide more accurate and scalable solutions for power systems [82].
+4 Proposed implementations for GridFM
+Figure 4: Applications: Examples of power system problems to be solved with GridFMs.
+The broader vision of a GridFM is the enablement of all or as many applications as possible.
+However, even if initially GridFMs would only be suitable for a single application, the superior adaptability of FMs over bespoke AI models through fine-tuning to thousands of different utilities with their specific data and requirements justifies the FM approach.
+In the above context, fine-tuning can mean different things.
+A GridFM may be fine-tuned to a set of diverse applications like load forecasting, fault detection and others.
+Initially, this may be too ambitious and challenging, considering the diversity of downstream applications as shown in Figure 4, each requiring different data, temporal scales, and pre-training strategies.
+But fine-tuning may also refer to tailoring a GridFM for a given application with user-specific (and potentially private) data to the particular user needs and requirements.
+For example, just in the United States, there are 3,200 utilities [71], with their own private and siloed data and grid-specific requirements.
+These utilities may not have sufficient data by themselves but could directly profit from the proposed approach.
+This multi-level fine-tuning approach is not unlike what can be observed in geospatial or weather foundation models, where FMs have shown excellent adaptability to different geographic regions, spatial and temporal resolutions for a given downstream application [72, 73, 50].
+We thus propose an implementation road map as outlined in Figure 5, with a first, more specific version of GridFMs, pre-trained for power flow applications (GridFM–v0), which is expected to provide real-world benefits in the next 1-3 year.
+In the long-term, we envision expanding and addressing more grid applications by including more diverse data modalities to pre-train a single GridFM or by combining different GridFMs, as discussed in Section 4.2.
+The rationale to start with power flow is as follows:
+(1) A highly relevant pre-training task
+FMs are known to excel when the pre-training task is closely related to the enabled downstream applications [74].
+Consequently, with AC power flow at the core of power system analysis [75], its reconstruction is an ideal pre-training task for GridFM–v0.
+This enables power flow related applications in the area of load flow (optimal power flow, state estimation, power flow), system security (contingency analysis, cascading failure), and beyond.
+(2) AI power flow models are computationally more efficient
+Current power flow solvers, e.g.
+using Newton-Raphson methods, are computationally expensive, especially for large grids.
+For instance, for an N-1 contingency analysis on the Western Interconnection, it takes approximately 5 minutes per scenario [76], while meeting the NERC TOP-001-3 requires running these analyses within a limited time period.
+In contrast, GridFM–v0 is projected to achieve a computational speed-up of 3 to 4 orders of magnitude over conventional solvers, as demonstrated on IEEE grid cases with up to 118 buses [77].
+For larger grids, we anticipate a similar or greater speedup, as GNNs exhibit linear complexity in the number of grid buses [78], while Newton-Raphson methods scale quadratically with the number of buses [79].
+This significant efficiency gain will enable real-time analysis of multiple scenarios of more complex grids [77], which will become crucial for maintaining system stability and optimizing future grids under uncertainties stemming from the energy transition and from climate change [76].
+(3) There have been significant advances in GNN-based power flow
+Using GNNs to represent the power grid and mapping power flow to the message-passing scheme of GNNs has been successfully demonstrated recently [77].
+These models replicated the results of traditional power flow solver with less than 1% error on the predicted quantities, despite past GNN training challenges [80].
+While these models were trained in a supervised way, recent results of the previously discussed GNN-based weather FMs [72, 73, 50] indicate that learning power flow with GNNs in a self-supervised way is a valid approach for a GridFM.
+Furthermore, large-scale data foundation models (FM) based on GNNs have already been developed [48], and a comprehensive review can be found in [81].
+Figure 5: GridFM implementation road map with four near-term phases to develop GridFM–v0, a GridFM for power flow related applications.
+The icons refer to academia, power system industry, and AI community.
+The long-term goal is creating a family of GridFMs for a broader range of grid challenges.
+4.1 Near-term: GridFM–v0, a GridFM for power flow applications
+The four phases as foreseen for the development of GridFM–v0, and as depicted in Figure 5 are discussed in detail in the following.
+4.1.1 Phase 1: Identify, collect and generate training data
+Training GridFM–v0 requires a large collection of solved power flow problems for different grid topologies, parameters, and different load conditions.
+In recent years, large power grid data-sets have been made available, created by (1) collecting grid topology and measurement data, (2) generating varying synthetic load conditions per grid model and (3) using established solvers to compute a wide range of power flow solutions under different operational conditions.We aim at using the datas-et provided by [13], which contains power flow solutions under real load conditions on 4 different grid topologies.
+Additionally, to complement this data, diverse power flow problems using traditional power flow solvers [82, 83, 84] will be computed, with power grid topologies from PGLIB-OPF [85] and standard IEEE benchmarks and with topology and load perturbations.
+Consequently, our model will be pre-trained on data from realistic topologies with a broad range of operational conditions.
+This approach facilitates the model to generalize well across topologies, as demonstrated in [86].
+For model fine-tuning, data-sets from [86] for system security analyses and [87] for Optimal Power Flow (OPF) will be used.
+The latter contains 300,000 solved optimal power flow problems for ten different grid topologies each, including solutions for N-1 contingency analysis, with grid sizes ranging from 14 to 14,000 buses.
+4.1.2 Phase 2: Model architecture development
+Figure 6: GridFM–v0 is pre-trained to reconstruct masked power flow data from input graphs.
+In the second phase, suitable architectures will be evaluated by pre-training performance analysis, and by adjusting models, training strategies, and loss functions.
+Figure 6 illustrates the Masked Auto-Encoder (MAE) pre-training task of GridFM–v0.
+Intuitively, the grid is modeled as a graph, where nodes represent electrical buses with active power, reactive power, voltage magnitude, and voltage angle, $(p_{i},q_{i},v_{i},\delta_{i})$, as node variables.
+We opted for this graph representation of power systems, but others can also be considered [88].
+Transmission lines and transformers are modeled as graph edges.
+An autoencoder [89, 90] is trained in a self-supervised manner to reconstruct masked node features from power flow solutions, each corresponding to a specific topology and load condition.
+This amounts to solving the power flow equations.
+During training, the loss function ensures meaningful reconstruction of node variables within the physical constraints.
+The total loss $\mathcal{L}=\mathcal{L}_{SCE}+\mathcal{L}_{powerflow}$ combines the Scaled Cosine Error (SCE) [89], which minimizes the distance between the original and reconstructed graphs, and $\mathcal{L}_{powerflow}$, which ensures that the power flow equations are satisfied.
+Recent studies show that an approach using physics-informed methods is the most effective for solving OPF.
+For example, [91] demonstrated that a hybrid loss function provides high-fidelity approximations for large-scale OPF, while [92] confirmed that this approach best solves AC OPF by integrating data-driven learning with physical laws.
+FMs, as a concept, are designed to incorporate diverse methods, including GNNs and PINNs, into a unified framework, enabling them to generalize effectively and adapt to various tasks, such as solving OPF problems.
+4.1.3 Phase 3: Model scaling and validation on industry data
+Due to regulatory constraints, it is unlikely or at least challenging to gain full access to sensitive data from operational power grids [93, 94, 95].
+To ensure the model’s accuracy and relevance in practical applications, it is therefore imperative that a GridFM be validated on operational data and real grid topologies.
+To address these challenges, an iterative approach is adopted.
+Our GridFM–v0 will initially be pre-trained and tested on openly available data.
+Subsequently, it will be re-trained and validated in collaboration with electricity utility partners behind the firewall in a secure compute environment.
+The results of these validations will guide us to generate more suitable synthetic training data, update the architecture, loss functions etc.
+so that the improved GridFM yields satisfactory results in real-world conditions once fine-tuned.
+This approach ensures that the GridFM remains performant and aligned with industry needs without exposing nor pre-training the model on proprietary data.
+4.1.4 Phase 4: Implement power flow-based use cases
+GridFMs are anticipated to achieve a computational speed-up of 3 to 4 orders of magnitude over traditional power flow solvers, as stated previously, which will directly benefit various applications:
+(1) Complementing Numerical Power Flow Solvers
+For applications that require solving power flow frequently, and where speed is critical, linear approximations (such as DC power flow [96]) are typically used.
+For these applications, faster AI-based power flow solvers offer a trade-off between “traditional” DC power flow and AC power flow solutions.
+We expect minimal fine-tuning of these models, as the task is equivalent to the suggested pre-training reconstructing task.
+(2) Contingency Analysis
+Contingency analysis is essential for assessing grid resilience.
+On a grid with 1000 lines, simulating the loss of each individual line requires solving 1,000 distinct power flow problems.
+However, when considering scenarios that involve the simultaneous loss of any two components, the number of power flow analyses increases dramatically to $\binom{1000}{2}=499,500$.
+By utilizing a power flow solver that is 1000 times faster, operators will be able to perform contingency analyses across a much larger set of scenarios.
+For example, solving 499,500 scenarios with AI-based methods would take roughly the same time as solving only about 500 scenarios using conventional physics-based solvers.
+Therefore, a GridFM–v0 will greatly enhance grid analyses and contribute to operating and planning under uncertainty.
+The power grid is continuously evolving, therefore, the GridFM–v0 must be able to account for different network topologies and different operating conditions of a same grid.
+This can be accomplished by training the model on grids with topological perturbations, as already demonstrated by [97], or by additionally training the same model on entirely different grid topologies, as done in [86].
+(3) Optimal Power Flow (OPF)
+The authors in [97] trained a model on synthetic data and demonstrated an OPF speedup of 2 to 4 orders of magnitude while maintaining errors of less than 1% in the objective value.
+Solving the OPF translates into fine-tuning the GridFM for reconstructing the generator power to minimize generation costs.
+Reconstructed node variables must adhere to bounds (e.g., generator capacity constraints), making this problem challenging.
+However, GridFM–v0 can be hierarchically integrated into optimization schemes where generator power is iteratively adjusted, solution feasibility is checked, and other variables are determined using the power flow solver.
+4.2 Long-term Vision: GridFMs for a broader set of applications in power systems
+In Section 4.1, we explored an initial version of GridFM for power flow applications, already addressing important challenges of the industry.
+The long-term objective, however, is to address as many applications from Figure 4 as possible, using a single GridFM or a family thereof, fully exploiting FM capabilities.
+Here we discuss how to build new industry-relevant applications on top of GridFM–v0.
+4.2.1 Extending GridFM–v0 to a GridFM capable of handling new applications
 Forecasting
-Forecasting is used in power grids for multiple tasks, including demand prediction, renewable production forecasting, and electricity price prediction.
-Demand and renewable forecasting are vital in ensuring power generation-demand balance, which is becoming increasingly important in a grid with high renewable penetration.
-The self-supervised learning paradigm used in FMs and their predictive ability make them intrinsically suitable for forecasting [85].
-Moreover, FMs can integrate exogenous variables, such as numerical weather predictions, which also carry associated forecast uncertainties.
+GridFM–v0 was pre-trained to reconstruct power flow data for a single time step.
+This can be readily extended to pre-training to reconstruct power data over multiple consecutive time steps, similar to Weather FMs [73].
+This enhancement will enable the forecasting of future grid states.
+By incorporating weather and climate data into the node embeddings, we will further improve the model’s ability to predict generation and anticipate weather-induced grid disturbances.
+Transient and Dynamic Stability Analysis
+In addition to training GridFMs on static power flow data, like in the case of GridFM–v0, the pre-training can additionally involve reconstructing high-resolution real-time data obtained from PMUs, which enables effective transient state estimation.
+This approach will facilitate rapid fault current detection and localization, efficient system restoration, and disturbance mitigation [98].
+Additionally, the generator rotor angle and voltage differential-algebraic equations can be added to the model using physics-based machine learning techniques [99] to, e.g., assess the grid’s stability margins.
 Control operations
-Here, the concerns include frequency and voltage control, network reconfiguration management, protection activation, and load shedding.
-FMs will help optimize and control the power system through sequential decision-making agents [86].
-A GridFM can enhance reinforcement learning (RL) agents by introducing prior knowledge of the environment before and during training, improving learning efficiency, and reducing risks.
-Electricity markets
-Electricity markets will benefit from the computational efficiency and predictive ability of a GridFM.
-The different markets, including futures, balancing, day-ahead, and intraday markets, can be optimized using the base GridFM or fine-tuned versions.
-Furthermore, the model will identify the security-based re-dispatch after the day-ahead market clearing in a timely manner [87].
-It will also help to pinpoint the location, amount, and type of ancillary services from distributed sources for congestion management in the future power grid with high shares of renewables.
-Finally, a GridFM will provide prompt solutions for economic dispatch and unit commitment problems, traditionally solved as linear and mixed-integer optimization problems [88].
-System security
-ML methods have been used to identify faulty components [12] and to speed up simulations like $N-1$, $N-K$ contingency, and cascading failure analyses [89], all critical tasks for operation and planning.
-Such models are faster and better leverage past knowledge for security predictions than traditional tools.
-Nevertheless, these analyses require additional encoding to incorporate topology with bus or branch-level features to inform about the final state of the entire grid.
-In this context, a GridFM tuned for system security will explore a broader and more complete set of scenarios, and thus both aid in designing a reliable power system and provide a system security tool for power system operations planning.
+There are two concrete ways in which GridFMs will help in control operations.
+First, the generative capabilities of FMs as a conditional load and generation scenario generator, can be used to construct scenario trees that are inputted to multistage stochastic programming algorithms, as done in [100, 101].
+Second, we can use GridFM–v0 and GridFMs in general as digital twins of power grids within Reinforcement Learning (RL) environments.
+Leveraging digital twins (data-driven representations of systems) enables faster computation of state transitions during RL training.
+This also allows RL agents to be trained on systems with highly complex dynamics that are difficult to model explicitly, such as distribution networks.
+As demonstrated in [102], a digital twin of a distribution network was used to train an RL agent, improving decision-making in both planning and operations.
+Third, GridFMs can also be used as a pre-trained representation for RL agents, for which FMs have been shown to be a valuable asset [103].
 Expansion planning
-Rapid electrification of the industrial and automotive sectors and climate change’s impact on generation profiles make expansion planning a crucial task at distribution and transmission levels.
-The generative capabilities of a GridFM to create future scenarios for load and generation (e.g., using public projections of electric vehicle sales and national renewable energy plans) will enhance decision-making.
-By integrating scenario generation with graph generation and deep reinforcement learning techniques, optimized and upgraded power systems can be designed to capture a broader spectrum of future extreme scenarios [90].
-Additionally, by exploiting cross-country information (e.g., the impact of interconnection flows), a GridFM will improve expansion planning strategies to prepare for these effects and enable the energy transition in the following decades.
+The pre-training of GridFM–v0, which involves reconstructing node-level masked power flow data, can be complemented by edge masking and reconstruction.
+Extending this approach further, we will leverage the model’s generative capabilities to also propose new grid topologies that optimize for capacity expansion.
+By integrating constraints such as network reliability, load growth forecasts, and economic cost factors, GridFM can generate topologies that minimize congestion, maximize redundancy, and ensure efficient power distribution.
+Optimizing grid topology and resource dispatch co-optimization has already been achieved with GNNs in [104].
 Cybersecurity
-Cybersecurity challenges could be tackled by leveraging FMs to speed up power system simulations.
-These FMs facilitate risk assessment and incident response in transmission and distribution systems by identifying and prioritizing threat scenarios [91] and designing optimal responses to maintain grid reliability [92].
-They can generate synthetic cybersecurity data by learning complex power system dynamics to overcome data scarcity, improving intrusion detection, and enabling dynamic stress testing against intelligent adversaries [93].
-A cyber-physical FM can combine SCADA and AMI data with cybersecurity information to enhance cyber threat mitigation, detection, and response.
-5 Conclusion
+GridFMs can be fine-tuned for anomaly detection tasks to identify patterns that signal potential cyber intrusions or physical failures [105].
+We can also use GridFMs to identify vulnerabilities and conditions that lead to cascading failures [106].
+Although the use cases mentioned above are the most apparent, experience with FMs and LLMs shows that once a model is available, numerous new applications can emerge.
+Therefore, the reader is encouraged to begin considering these possibilities, as FMs for the power grid will soon become a reality.
+4.2.2 Building multi-modal GridFMs
+Ideally, GridFMs should be able to handle more than one set of application tasks.
+This can be done by combining multiple GridFMs together [103] or by building self-contained stand-alone larger models, ingesting multi-modal spatial, temporal, as well as text data [14].
+Finally, concepts such as Mixture-of-Experts (MoE) frameworks [107] can be exploited to compose GridFMs from a family of GridFMs.
+Regardless of the method used, GridFMs will need to handle diverse data modalities in the long term, to account for complex external factors—such as climate, weather, economic and societal influences, distributed power sources, and the electrification of mobility and heating [76].
+Training multi-modal FMs requires creating tokens for each modality, aligning them in space and time, and potentially normalizing or reducing their dimensionality [108].
+Consequently, tasks beyond power flow and grid topology reconstruction [109] will need to be incorporated for pre-training.
+This approach has been demonstrated with models like 4M [110], which unifies different modalities into discrete tokens for computer vision.
+We anticipate that the number of trainable parameters for such models could approach those of weather models, reaching hundreds of millions [56, 111].
+5 Conclusion and Discussion
 The rapid emergence of FMs is reshaping the AI landscape.
-While such models are well established for text and image analysis, here we have introduced the idea of a GridFM: an FM for the power grid, able to learn from time-series, geospatial, and other data to improve power grid planning, operations, and control.
-We first motivated the need for a GridFM by pointing to the growing gaps between computational capabilities and the industry’s need to cope with the increasing uncertainty and complexity of the power grid, largely due to the energy transition.
-We then reviewed the three core advantages of the FM approach for power systems: excellent missing token prediction skills (signal reconstruction); broad adaptability and scalability; and simulation acceleration.
-In another domain characterized by apparently intractable complexity, namely weather forecasting, we have seen AI weather FMs approximate the fundamental equations governing weather to produce much more computationally efficient models (more than 4 orders of magnitude speed-up).
-In a similar manner, we propose to develop a GridFM that learns the behaviors and dynamics from measurements and simulations for different network configurations.
-The superior computational performance of such a GridFM has the potential to accelerate grid analyses by several orders of magnitudes.
-This acceleration will, in turn, enable modeling of the increasingly complex (e.g., bidirectional power flow, combined transmission, and distribution) and uncertain grid (e.g., by modeling different scenarios going beyond the traditional N-1 analysis).
-The promise of such a GridFM is that once pre-trained on multiple grids and various networks, it can be readily fine-tuned to specific grid topologies (distribution and transmission).
-In summary, we argue that research in the proposed FM for the power grid is imperative to facilitate the energy transition.
-We envision the power system and AI communities collaborating to take on this “moonshot” and thus revolutionize power grid analysis, leveraging AI advancements to meet common goals and needs.
-Acknowledgement
+While FMs are well established for text and image analysis, here we have introduced the idea of GridFMs: FMs for the power grid, able to learn from grid-relevant data to improve power grid operations, planning, and control.
+We first highlighted the strengths of FMs, which include excellent token prediction skills (signal reconstruction and forecasting), broad adaptability and scalability, and simulation acceleration.
+To demonstrate that GridFMs are well suited to close the growing gaps between computational capabilities and the industry’s need to cope with the increasing uncertainty and complexity of the power grid, we draw on the experiences of FMs for weather forecasting.
+The latter is a domain characterized by apparently intractable complexity.
+Nevertheless, we have seen scalable and adaptable FMs to produce much more computationally efficient models by more than 4 orders of magnitude than traditional weather forecasting models.
+We have outlined concrete implementation phases to create an initial version, GridFM–v0, which is pre-trained on power flow data and tailored to accelerate power flow-related downstream applications (e.g., optimal power flow and contingency analysis).
+In the same way that FMs have been able to learn the fundamental equations governing weather, we envision GridFM–v0 to learn power flow, thereby creating a much more efficient model.
+The key advantage of such a GridFM–v0 is that once pre-trained on multiple grids and various networks and operating conditions, it can be readily fine-tuned to specific grid topologies and user requirements based on their own private data.
+This paper advocates for research and development of FMs for the power grid, to facilitate the energy transition through collaboration between the various stakeholders.
+All stakeholders will benefit from GridFMs, simply by fine-tuning it to their specific needs and with their proprietary data and knowledge – without re-developing a full bespoke AI model each time.
+Towards that end, we envision, under the leadership of industry associations and open source initiatives, for the community of stakeholders to join forces and make the required investments to research and develop GridFMs in open source, thereby revolutionizing power grid analysis and leveraging AI advancements to meet common goals and needs.
+Author contributions
+  • 1.
+    H.F.H.: Conceptualization, Investigation, Writing – Original Draft, Project Administration
+  • 2.
+    B.G., L.S.A.M., A.P., A.V., J.W.: Conceptualization, Investigation, Writing – Original Draft
+  • 3.
+    J.B., A.B.M., T.B., S.C., I.F., B.H., R.J., K.K., V.M., F.M., M.D.M., O.R., H.S., L.X., E.S.Y., A.Z.: Investigation, Writing – Original Draft
+  • 4.
+    A.J.B., R.J.B., B.P.B., J.S., S.S.: Writing – Review & Editing
+Lead contact
+Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Hendrik F. Hamann (hehamann.hh@gmail.com).
+Acknowledgment
 The work was supported by the respective organizations of the authors.
-The authors acknowledge the invaluable discussions and interactions in the working group for Foundation Models for the Power Grid, including Roya Ahmadi, Rich Argentieri, Gianni Vittorio Armani, Venkat Balachandran, Ranjini Bangalore, Venkat Banunarayanan, Edmon Begoli, Larry Bekkedahl, Keith Benes, Michelle Berti, Budhu Bhaduri, Roberto Bianchessi, Gilbert Bindewald, Louise Blais, David Bromberg, Matthias Bucher, Eren Cam, Supriya Chinthavali, Seong Choi, Julien Chosson, Joseph Collum, Magnus Cormack, Peter De Bock, Fei Ding, Kathy Duviella, Darlain Edeme, Yamshid Farhat, Mark Fenton, Marcus Freitag, Alirez Ghassemian, Tassos Golnas, Gavin Goodland, Luca Grella, John Grosh, Benois Grossin, Andy Hackett, Hagen Haentsch, Zeryai Hagos, Grahame Hay, Frank Hellmann, Andy Henton, Raffael Hilber, Dan Hilton, Gabriela Hug, Alex Hurley, Sorana Ionescu, Kevin Johnston, Gareth Jones, Adrian Kelly, Bernd Klockl, Teddy Kott, Benjamin Kroposki, Teja Kuruganti, Larry Light, Dalton Lunga, Matt Magill, Adrian Maldonado, Antoine Maret, Pete Marsh, Russel Mason, James McClean, John Messer, Nikki Militello, Jordan Murkin, Xabi Muruaga, Indu Nambiar, Christian Nauck, Caleb Northrop, Gabby Nyirjesy , Marco Padula, Patrick Panciatici, Raul Alfaro Pelico, Mark Petri, Pierre Pinson, Abhishek Pontis, Josh Power, Alan Pressman, Feng Qiu, Mark Rafferty, Pippa Robertson, Nicoletta Rocca, Caroline Rose-Newport, Lyndon Ruff, Bryan Sacks, Kelly Sanders, Giovanni Sansavini, Michael Sheppard, Robert Shorten, Vikas Singhvi, Stanislav Sobolevsky, Pratik Sonthalia, Brian Spears, Declan Stock, Olle Sundstroem, Matthew Tarduogno, Roger Thompson, Anna Carolina Tortora, Alex Vivas, Lucy Vu, Ben Weiner, Casey Werth, Daniel Wetzel, Gary White, Xin Wu, Steven Wysmuller, Yunchi Yang, Thomas Zadlo, Marzia Zafar and many others.
+The authors acknowledge the invaluable discussions and interactions in the working group for Foundation Models for the Power Grid, including Roya Ahmadi, Rich Argentieri, Gianni Vittorio Armani, Venkat Balachandran, Ranjini Bangalore, Venkat Banunarayanan, Edmon Begoli, Larry Bekkedahl, Keith Benes, Michelle Berti, Budhu Bhaduri, Roberto Bianchessi, Gilbert Bindewald, Louise Blais, David Bromberg, Matthias Bucher, Eren Çam, Supriya Chinthavali, Julien Chosson, Joseph Collum, Magnus Cormack, Peter De Bock, Fei Ding, Kathy Duviella, Darlain Edeme, Yamshid Farhat, Mark Fenton, Marcus Freitag, Alirez Ghassemian, Tassos Golnas, Gavin Goodland, Luca Grella, John Grosh, Benois Grossin, Andy Hackett, Hagen Haentsch, Zeryai Hagos, Grahame Hay, Frank Hellmann, Andy Henton, Raffael Hilber, Dan Hilton, Gabriela Hug, Alex Hurley, Sorana Ionescu, Kevin Johnston, Gareth Jones, Adrian Kelly, Bernd Klockl, Teddy Kott, Benjamin Kroposki, Teja Kuruganti, Larry Light, Dalton Lunga, Matt Magill, Adrian Maldonado, Antoine Maret, Pete Marsh, Russel Mason, James McClean, John Messer, Nikki Militello, Jordan Murkin, Xabi Muruaga, Indu Nambiar, Christian Nauck, Caleb Northrop, Gabby Nyirjesy , Marco Padula, Patrick Panciatici, Raul Alfaro Pelico, Mark Petri, Pierre Pinson, Abhishek Pontis, Josh Power, Alan Pressman, Feng Qiu, Mark Rafferty, Pippa Robertson, Nicoletta Rocca, Caroline Rose-Newport, Lyndon Ruff, Bryan Sacks, Kelly Sanders, Giovanni Sansavini, Michael Sheppard, Robert Shorten, Vikas Singhvi, Pratik Sonthalia, Brian Spears, Declan Stock, Olle Sundstroem, Matthew Tarduogno, Roger Thompson, Anna Carolina Tortora, Alex Vivas, Lucy Vu, Ben Weiner, Casey Werth, Daniel Wetzel, Gary White, Xin Wu, Steven Wysmuller, Yunchi Yang, Thomas Zadlo, Marzia Zafar and many others.
 This material is based upon work supported by the U.S.
 Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.
-Author contributions
-Conceptualization, H.F.H., L.S.A.M.
-and J.W.; Investigation, H.F.H., T.B., B.G., L.S.A.M., A.P., A.V., J.W., J.B., A.B.M., S.C., I.F., B.H., R.J., K.K., V.M., F.M., M.D.M., O.R., H.S., L.X., E.S.Y., and A.Z.; Writing – Original Draft, H.F.H., T.B., B.G., L.S.A.M., A.P., A.V., J.W., J.B., A.B.M., S.C., I.F., B.H., R.J., K.K., V.M., F.M., M.D.M., O.R., H.S., L.X., E.S.Y., and A.Z.; Writing – Review & Editing, A.J.B., R.J.B., B.P.B., J.S.
-and S.S.; Project Administration, H.F.H.
 Declaration of Interests
 The authors declare no competing interests. H.F.H is also affiliated with Yamagata University and University of Illinois Urbana-Champaign. S.S. is also affiliated with Masaryk University.
+Declaration of Generative AI and AI-Assisted Technologies
+During the preparation of this work, the authors used ChatGPT to find more concise reformulations and to condense paragraphs.
+After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.